The combined SAT scores for the students at a local high school are normally distributed with a mean of 1517 and a standard deviation of 307. The local college includes a minimum score of 1179 in its admission requirements. What percentage of students from this school earn scores that satisfy the admission requirement?

P(X > 1179) = %

Enter your answer as a percent accurate to 1 decimal place (do not enter the "%" sign). Answers obtained using exact z-scores or z-scores rounded to 3 decimal places are accepted.
Answer #1

Solution :

Given that,

mean = = 1517

standard deviation = = 307

P(x >1179 ) = 1 - P(x<1179 )

= 1 - P[(x -) / < (1179-1517) / 307]

= 1 - P(z < -1.101)

Using z table

= 1 -  0.1354

=0.8646

answer=86.5